OpenStudy (anonymous):

Differential Equations help?? Find two different solutions to initial value problem, why there is no contradiction of the Existence Uniqueness Theorem? dy/dx=2x(y-5)^(3/4) y(1)=5 Differential Equations help?? Find two different solutions to initial value problem, why there is no contradiction of the Existence Uniqueness Theorem? dy/dx=2x(y-5)^(3/4) y(1)=5 @Mathematics

OpenStudy (turingtest):

I can solve it but I can't see how to get two answers...

OpenStudy (anonymous):

I don't see how you can ever get two solutions with these types of questions

OpenStudy (anonymous):

like another one is y'=y^(1/5)

OpenStudy (anonymous):

unless it's like ln and you can use absolute?

OpenStudy (turingtest):

Well I'm looking at it... James knows though.

OpenStudy (turingtest):

The other one y'=y^(1/5) is not an IVP without the initial value though, so it makes sense it has different answers depending on the value of C you get.

OpenStudy (jamesj):

Hm. If y'=2x(y-5)^(3/4) then \[ (y-5)^{-3/4} dy = 2x dx \] \[ \implies (1/4)(y-5)^{1/4} = x^2 + C \] \[ \implies y - 5 = [ 4(x^2 + C) ]^4 \] \[ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ = 4^4 (x^2 + C)^4 \] With the initial condition y(1) = 5, it must be that C = -1 hence \[ y = 2^8 (x^2 - 1)^4 \] So I guess I don't see how there are two solutions.

OpenStudy (jamesj):

sorry ... final solution is \[ y = 2^8 (x^2 -1)^4 + 5 \]

OpenStudy (anonymous):

second initial is y(0)=0

OpenStudy (anonymous):

I think it must exist since it ask why does these 2 solutions not contradict existence uniqueness

OpenStudy (turingtest):

Well there you have it I guess. I can say James is usually right about this kind of thing.

OpenStudy (jamesj):

Oh, I see. The uniqueness theorem says that if f'(x) = g(x,y) and g is reasonably behaved, then given an initial condition y(a) = b, then the equation has a unique solution. So saying the equation \( y'=2x(y-5)^{3/4} \) has a solution for one initial condition and a different solution for a different initial condition doesn't violate the uniqueness theorem, because ...

OpenStudy (jamesj):

the hypotheses of the uniqueness theorem aren't satisfied. Instead, you're showing equation + initial condition A => solution A; and equation + initial condition B => solution B and these two solution aren't the same. Great, no problem.

OpenStudy (jamesj):

What the unique theorem says is that if equation + initial condition A => solution A equation + initial condition A => solution B then solution A = solution B

OpenStudy (turingtest):

So we make up another not-given initial condition to get our other equation?

OpenStudy (anonymous):

Oh I see so if you have two solutions then they are not related but if they are equal then the initial conditions should be the same

OpenStudy (jamesj):

Yes. Given the initial conditions y(1) = 5, we get one solution (above) y(0) = 0, we will get another solution (not yet written out)

OpenStudy (turingtest):

Sweet, thanks.

OpenStudy (anonymous):

I don't see how they are two different solutions though? I get that they are two different initial conditions

OpenStudy (anonymous):

but they have the same value, oh so that goes along with the Existence Uniqueness theorem

OpenStudy (jamesj):

By the way, it can be the case that equation + initial condition A => solution S equation + initial condition B => solution S That's not a contradiction.

OpenStudy (jamesj):

your equation + initial condition y(1) = 5 => \( y = 2^8 (x^2 - 1)^4 + 5 \) That's a solution and indeed THE solution given the ODE and the IC. Now with the ODE and the IC y(0) = 0 you will have a different solution. I haven't written it down. You write it down.

OpenStudy (anonymous):

Yeah I did they both are 0

OpenStudy (anonymous):

but yeah I get it now, they don't contradict b/c of the condition just wrote up top

OpenStudy (jamesj):

So write down the function y(x) for me for y(0) = 0

OpenStudy (anonymous):

Isn't it the same but with a different initial condition? Or am I doing this wrong. I just plugged in y(0)=0

OpenStudy (jamesj):

And so what then is the solution of the ODE: dy/dx=2x(y-5)^(3/4) + IC: y(0) = 0

OpenStudy (jamesj):

what function satisfies that ODE and IC.

OpenStudy (anonymous):

dy/dx=0? ==> 2(0)(0-5)^(3/4)=0

OpenStudy (jamesj):

The function that satisfies ODE: dy/dx=2x(y-5)^(3/4) + IC: y(1) = 5 is y(x) = 2^8 (x^2 - 1)^4 + 5

OpenStudy (jamesj):

No.

OpenStudy (anonymous):

Ohhhh I get it you have to integrate

OpenStudy (anonymous):

that function is for both initial condition right?

OpenStudy (jamesj):

I've already solved the equation in general: look above. The procedure is to take that general solution and then substitute the initial condition into that general solution to solve for the constant C introduced in the general solution.

OpenStudy (jamesj):

Namely, the general solution of the ODE dy/dx=2x(y-5)^(3/4) is \[ y(x) = 2^8 (x^2 + C)^4 + 5 \]

OpenStudy (jamesj):

Now, for the IC y(1) = 5, I substituted in that condition to solve for C. That gave me THE solution the Initial Value Problem --- ODE + IC --- of \( y(x) = 2^8 (x^2 -1)^4 + 5 \).

OpenStudy (anonymous):

Oh okay I gotcha then you sub in y(0)=0 and find C for that case

OpenStudy (jamesj):

Yes.

OpenStudy (anonymous):

and that gives two equations/solutions. Okay thanks for your help! Makes a lot of sense now

OpenStudy (jamesj):

Actually, it looks to me as though there is no solution for y(0) = 0. Write \( dy/dx = g(x,y) \) where \( g(x,y) = 2x(y-5)^{3/4} \). What's the domain of g? \[ Dom(g) = \mathbb{R} \times [5,\infty) \] Hence (0,0) isn't in the domain of g and we shouldn't expect to find a solution to the ODE with IC y(0) = 0.

OpenStudy (anonymous):

Then how do you know what initial condition to use for the second solution? Aren't you trying to get an initial where dy/dx=2x(y-5)^(3/4) is equal to zero?

OpenStudy (anonymous):

can I do like y(0)=5? should work as long as y> = 5